OpenStudy (anonymous):

What is the correct solution to the equation 4(x – 3) = 16 ?

OpenStudy (anonymous):

4(x-3)= 16 4x-4(3) = 16 4x - 12 = 16 solve for x

OpenStudy (anonymous):

4x-12=16-x+2 +12 +12 4x=30-x +x +x 5x=30 /5 /5 x=6

OpenStudy (anonymous):

You know that 4x4 equals 16. So, x-3 has to equal to 4. So, x=7 because 7-3=4.